OpenStudy (anonymous):

find the slope if it exists 7x-5y+19=0 answer should be an integer or simplified fraction unless it is undefined

OpenStudy (saifoo.khan):

Slope = 7/5

OpenStudy (chriss):

7x-5y=-19 -5y=-7x-19 \[y=\frac{-7}{-5}x-\frac{19}{5}\] y=mx+b m=slope \[slope=\frac{7}{5}\]

OpenStudy (anonymous):

slope = 1,4